OpenStudy (anonymous):

Find f(-3) if f(x) = x 2. -9 -6 6 9

OpenStudy (anonymous):

\[f(-3)=(-3)^2=???\]

OpenStudy (anonymous):

assuming that \[f(x)=x^2\]

OpenStudy (anonymous):

oh! -6?

OpenStudy (anonymous):

\[(-3)*(-3)=(-3)^2\] it is not -6

OpenStudy (anonymous):

oh...9! I added haha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yay! haha, could you help me figure out this: Find x if f(x) = 2x + 7 and f(x) = -1.

OpenStudy (anonymous):

there is no question there

OpenStudy (anonymous):

Yes there is lol its "Find x if f(x) = 2x + 7 and f(x) = -1. "

OpenStudy (anonymous):

oh misread they're both equal to f(x) so 2x+7=-1 right?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

add 7 to -1, or multiply 7 to -1?

OpenStudy (anonymous):

subtract 7 on both sides to get rid of the 7 on the left side

OpenStudy (anonymous):

oh ok....it comes out in all to be -3. did I do that right?

OpenStudy (anonymous):

no you did not

OpenStudy (anonymous):

darn it!! well I subtracted and got 6

OpenStudy (anonymous):

\[2x+7=-1\] \[2x+7-7=-1-7\] \[2x=-1-7\] -1-7 does not equal 6

OpenStudy (anonymous):

oh -8